# Driving as a Diagnostic Tool: Scenario-based Cognitive Assessment in Older Drivers from Driving Video

### v2

@@ 4.2 Percentage of Misclassification
Figure 8: Percentage of misclassification vs. - (a) COGSTAT score and (b) MoCA score. The relevant route segments were filtered out from the Omaha most transited route database shown in Figure 3.
@@
Figure 8 displays the relationship between misclassification percentage and cognitive scores. An overall trend observed was that subjects labeled as MD-aging tend to exhibit higher misclassification percentages compared to the Normal-aging drivers. This is evident in the middle cognitive score ranges (COGSTAT: 580–650; MoCA: 12–20), where the misclassification is higher. It suggests that the middle zones correspond to a transition region where the classifier shows more uncertainty, due to the overlapping feature space between early-stage impairment and Normal-aging. In addition, it underscores the anticipated challenge of reliably distinguishing subtle cognitive impairments from Normal-aging. On the other hand, subjects with low (COGSTAT <550<550 and MoCA <8<8) and high (COGSTAT >700>700 and MoCA >20>20) cognitive scores consistently yield lower misclassification rates. This behavior indicates that the classifier learned prominent feature representations in these regions and it is able to classify the aging groups more confidently. Additionally, it reinforces the validity of using COGSTAT and MoCA as effective biomarkers in well-separated diagnostic states. Overall, these trends are consistent with the model’s performance as reported in Table 3. The global performance and localized misclassification patterns support the interpretation that the model’s predictive success is not random but rather driven by the prominent vision signals yielded from a certain cognitive state. Subjects with well-defined cognitive profiles were more reliably classified. However, uncertainty increased in transitional cognitive states, reflecting appropriate model sensitivity to the early-stage classification. This behavior supports the validity of our approach and suggests that further improvement may require incorporating auxiliary clinical signals to resolve the transition zone uncertainty. Overall, this subject-level metric captures individual classification stability and uncertainty at the subject level.
@@ 4.3 Limitations and Future Works
The current dataset is constrained to a limited number of driving scenarios, which might not cover the full range of real-world driving complexity (e.g., night driving, inclement weather, heavy traffic). Future work can include a broader set of driving scenarios to generalize findings across real-world traffic and environmental conditions. The driver-level separation was limited to a 3-fold cross-validation with a small pool of test subjects. This can lead to variance in performance estimation and may not generalize well to other populations and road infrastructures. Recruiting more diverse drivers and conducting longitudinal assessments would strengthen generalization. Additionally, the number of dementia subjects in the study was relatively small ($N=9$). Including more dementia participants would guarantee the statistical validity of the findings and provide deeper insights into the driving cues associated with dementia. Moreover, the use of in-vehicle video recording raises potential concerns regarding privacy. Future implementations should address these with robust data management. Another notable limitation is the relatively limited availability of high-quality in-vehicle video data per cognitive group. The model’s misclassification rates, particularly in the transitional region, were influenced by this data scarcity. We believe that scaling up the dataset with higher-quality video samples for both cognitively normal and dementia groups will allow the method to extract more meaningful and generalizable features. This would be especially critical for improving video analysis accuracy in cases with confusing behavioral features.
@@ 5 Conclusion
This study highlights the potential of monitoring real-world driving behavior as a valuable tool for early cognitive health assessment in older adults. The results suggest that scenario-based analysis plays a significant role in video-based cognitive assessment. By segmenting driving video data into semantically distinct scenarios, our approach enables targeted driving behavior analysis under specific cognitive demands. In the scope of this work, freeway interchange scenarios yield significantly higher classification performance compared to interstate scenarios under driver-level separation. Therefore, certain driving scenes inherently capture richer behavioral variance, which large vision models can leverage to distinguish between cognitive aging driving patterns. This approach offers a low-cost, scalable alternative to traditional cognitive screening methods, enabling passive and non-intrusive cognitive monitoring. It has direct implications for the healthcare industry, where early detection of cognitive decline remains a costly and time-consuming process. Additionally, the ability to extract digital biomarkers from routine driving behavior transforms personal vehicles into affordable diagnostic tools. This reduces patient burden while supporting continuous and longitudinal health assessment. Moreover, it holds promise in providing an efficient screening tool for aging populations in early detection of cognitive decline, especially in areas with limited access to healthcare facilities.
@@ 6 Acknowledgements
This work was supported by the National Institutes of Health (NIH), the National Institute on Aging (NIA 5R01AG17177-18), and the University of Nebraska Medical Center Mind & Brain Health Labs. The views expressed in this paper are those of the authors alone and not the NIH or NIA. We thank our research team for coordinating this project. Due to personally identifying information (PII), the data are not available online. The authors declare no potential conflict of interest.
@@ Appendix A Study Recruitment Criteria
The inclusion and exclusion criteria used for participant recruitment are summarized in Table 4.
@@
(b) residing independently in Nebraska (b) Congestive heart failure
or nearby states (e.g., Iowa) (c) major psychiatric illness
(c) possession of a valid driver’s license (d) active vestibular disease
(d) met Nebraska visual acuity standards (>20/50>20/50 OU, corrected or uncorrected) (e) substance abuse within one year prior to enrollment
(e) proficiency in English (f) use of medications that could confound study outcome
(f) capability to provide informed consent (g) clinical or neurological conditions (e.g., stroke) that could interfere with the study outcome
@@ Appendix B Cognitive Tests used for COGSTAT score
Table 5: Cognitive Tests conducted for the COGSTAT composite score.
@@
Note: If all cognitive tests are the same as the mean, final COGSTAT will be $14\times 50=700$

### v3

@@ 4.2 Percentage of Misclassification
Figure 9: Percentage of misclassification vs. - (a) COGSTAT score and (b) MoCA score. The relevant route segments were filtered out from the OMR database shown in Fig. 4.
@@
Fig. 9 displays the relationship between misclassification percentage and cognitive scores. An overall trend observed was that subjects labeled as MD-aging tend to exhibit higher misclassification percentages compared to the Normal-aging drivers. This is evident in the middle cognitive score ranges (COGSTAT: 580–650; MoCA: 12–20), where the misclassification is higher. It suggests that the middle zones correspond to a transition region where the classifier shows more uncertainty, due to the overlapping feature space between early-stage impairment and Normal-aging. On the other hand, subjects with low (COGSTAT <550<550 and MoCA <8<8) and high (COGSTAT >700>700 and MoCA >20>20) cognitive scores consistently yield lower misclassification rates. This behavior indicates that the classifier learned prominent feature representations in these regions and it is able to classify the aging groups more confidently. Additionally, it reinforces the validity of using COGSTAT and MoCA as effective biomarkers in well-separated diagnostic states. Overall, these trends are consistent with the model’s performance as reported in Table 3.
4.3 Comparison with Benchmark Approaches
Table 4: Comparison with a recent benchmark tested on the RWRAD dataset
Method Feature(s) $\mathbf{a}$ $\mathbf{P}$ $\mathbf{R}$ $\mathbf{F1}$
Joshi et al. [35] Demo 61.86 70.77 63.89 67.15
Driving 55.93 64.71 61.11 62.86
Sleep 62.71 71.21 65.28 68.12
Demo + Driving 66.95 71.43 76.39 73.83
Driving + Sleep 63.56 71.01 68.06 69.50
All 68.64 73.97 75.00 74.48
Prediction 70.48 71.88 77.97 74.80
Ours RF-Fwy-int. 71.03 75.97 77.28 76.62
+DLS† 72.32 77.49 82.95 80.13
$\Delta$ +3.68 +3.52 +7.95 +5.65
Demo: demographics, Driving: driving performance, Sleep: sleep log, All: five distinct features including demographics, driving performance, activity space and physiological states, cognitive status prediction for next year using previous year data, RF-Fwy-int.: RF classifier trained on OMR Freeway interchange routes, DLS: Driver-level Separation, †used max value, $\Delta$: gain over the best performing model (All)
To benchmark our framework, we compared the model performance with a recent approach proposed by Joshi et al. [35] shown in Table 4, which includes demographics, driving features and sleep log for baseline cognitive status classification using XGBoost [30]. Their best-performing model achieved 68.64% classification accuracy $\mathbf{(a)}$ and 74.48% $\mathbf{F1}$ score on the RWRAD dataset. In contrast, our framework outperformed these results on the Freeway interchange scenario, achieving 71.03% accuracy $\mathbf{(a)}$ and 76.62% $\mathbf{F1}$ score under random sampling. Additionally, for the same scenario, the maximum accuracy $\mathbf{(a)}$ and $\mathbf{F1}$ score went up to 72.32% and 80.13%, respectively, with the driver-level separation (DLS) applied. These results demonstrate that scenario-based naturalistic driving video can serve as a scalable and non-invasive biomarker for cognitive aging.
Table 5: Comparison with the recent clinical ML and DL methods
Method Modality Dataset $\mathbf{a}$ $\mathbf{F1}$
JMMLRC [8] MRI, SNPs ADNI-1 [29] 58.40 68.30
MRN [75] sMRI ADNI-1,2 [29] 63.23 60.23
LDA-ELM [43] MRI, PET ADNI [29] 66.70 64.90
SSMFS [73] MRI, PET ADNI-1 [29] 67.80 60.42
XADLiME [50] sMRI ADNI-1,2 [29] 69.05 77.47
LA-GMF [71] sMRI ADNI [29] 70.33 67.22
RF-Fwy-int. (ours) Video RWRAD 71.03 76.62
+ DLS†(ours) Video RWRAD 72.32 80.13
$\Delta$ +1.99 +12.91
ADNI: Alzheimer’s Disease Neuroimaging Initiative, SNP: single nucleotide polymorphisms, sMRI: structural MRI, PET: Positron Emission Tomography, RF-Fwy-int.: RF classifier trained on OMR Freeway interchange routes, DLS: Driver-level Separation, †used max value, $\Delta$: gain over the LA-GMF model
We compared our framework with recent clinical Machine Learning and Deep Learning methods and summarized the results in Table 5. Most of these approaches heavily rely on neuroimaging modalities such as MRI and PET, which are costly, invasive, and require significant time and clinical expertise to analyze. In contrast, our scenario-based video framework provides a lightweight, non-invasive alternative and outperforms most clinical ML and DL methods accuracy range of 58–70%. The most competitive baseline, LA-GMF [71], reported 70.33% accuracy and 67.22% F1-score. Our framework reached 72.32% accuracy and 80.13% F1-score, showing improvements of $\sim$2% in accuracy and +12.91% in F1-score. Additionally, while clinical models often fail to capture how cognitive decline manifests in everyday activities, our method leveraged naturalistic driving behavior to capture functional patterns directly from real-world scenarios.
Another recent study by Danda et al. [16] found that amyloid-positive older adults exhibited higher average speeds, intermittent speed variability and increased acceleration during freeway on-ramp driving (similar to Fwy-int scenario), suggesting less adaptive driving. In clinical research, Amyloid is a protein that accumulates abnormally in the brain and forms plaques linked to the development of MD. Patients with cognitive impairment who test amyloid-positive are typically diagnosed as having MD, whereas amyloid-negative patients are generally excluded from an MD diagnosis [65, 7]. These findings confirm that Freeway ramp driving requires increased cognitive load involving behavioral reactions to rapidly changing traffic and abrupt speed shifts, which combine perception with motor control [16]. Additionally, they underscore our results of higher classification accuracy ($\mathbf{a}$ = 71.03%) and F1-score ($\mathbf{F1}$ = 0.6712) in Freeway interchange scenarios compared to the Interstate settings ($\mathbf{a}$ = 55.12%, $\mathbf{F1}$ = 0.5368). Although their approach combined driving patterns with physiological signals, it validates the relevance of using contrastive and challenging roadway scenarios in cognitive assessment tasks. Therefore, this study supports our findings that the Freeway interchange scenario provides more informative and strong behavioral cues that the large vision model can leverage to better separate cognitive aging groups.
@@ 4.4 Limitations and Future Works
The current dataset is constrained to a limited number of driving scenarios, which may not cover the full range of real-world driving complexity (e.g., night driving, inclement weather, heavy traffic). Future work can include a broader set of driving scenarios to generalize findings across real-world traffic and environmental conditions. The driver-level separation was limited to a 3-fold cross-validation with a small pool of test subjects. This can lead to variance in performance estimation and may not generalize well to other populations and road infrastructures. Recruiting more diverse drivers and conducting longitudinal assessments would strengthen the generalizability of the findings. Additionally, the number of dementia subjects in the study was relatively small ($N=9$). Including more dementia participants would guarantee the statistical validity of the findings and provide deeper insights into the driving cues associated with dementia. Moreover, the use of in-vehicle video recording raises potential concerns regarding privacy. Future implementations should address these with robust data management. Another notable limitation is the relatively limited availability of high-quality in-vehicle video data per cognitive group. The model’s misclassification rates, particularly in the transitional region, were influenced by the scarcity of this data. We believe that scaling up the dataset with higher-quality video samples for both cognitively normal and dementia groups will allow the method to extract more meaningful and generalizable features. This would be especially crucial for enhancing video analysis accuracy in cases with complex behavioral features.
@@ 5 Conclusion
This study highlights the potential of scenario-based driving behavior monitoring as a valuable diagnostic tool for early cognitive decline assessment in older adults. The results suggest that scenario-based analysis plays a significant role in video-based cognitive assessment. By segmenting driving video data into semantically distinct scenarios, our approach enables targeted analysis of driving behavior under specific cognitive demands. In the scope of this work, the freeway interchange scenario yielded significantly higher classification performance compared to the interstate scenario. Therefore, certain driving scenes inherently capture richer behavioral variance, which large vision models can leverage to distinguish between cognitive aging driving patterns. This approach offers a low-cost, scalable and effective alternative tool to traditional cognitive screening methods, enabling passive and non-intrusive cognitive monitoring. It has direct implications for the healthcare industry, where early detection of cognitive decline remains a costly and time-consuming process. Additionally, the ability to extract digital biomarkers from routine driving behavior transforms personal vehicles into affordable diagnostic tools. This reduces patient burden while supporting continuous and longitudinal cognitive health assessment and monitoring disease progression. Moreover, it holds promise as an efficient screening tool for aging populations, particularly in early detection of cognitive decline, especially in areas with limited access to healthcare facilities.
@@ 6 Acknowledgements
This work was supported by the National Institutes of Health (NIH), the National Institute on Aging (NIA 5R01AG17177-18), and the University of Nebraska Medical Center Mind & Brain Health Labs. The views expressed in this paper are those of the authors alone and not the NIH or NIA. We thank our research team for coordinating this project. Due to the presence of personally identifiable information (PII), the data are not available online.
@@ Appendix A Study Recruitment Criteria
The inclusion and exclusion criteria used for participant recruitment in the RWRAD dataset are summarized in Table 6.
@@
(b) Residing independently in Nebraska or nearby states (e.g., Iowa) (b) Congestive heart failure
(c) Possession of a valid driver’s license (c) Major psychiatric illness
(d) Meets Nebraska visual acuity standards (>20/50>20/50 OU, corrected or uncorrected) (d) Active vestibular disease
(e) Proficiency in English (e) Substance abuse within one year prior to enrollment
(f) Capability to provide informed consent (f) Use of medications that could confound study outcome
(g) Clinical or neurological conditions (e.g., stroke) that could interfere with the study outcome
@@ Appendix B Cognitive Tests used for COGSTAT score
Table 7: Cognitive tests for the COGSTAT composite score.
@@
Note: If all test scores == mean score, then the final score will be $14\times 50=700$